Crop Monitor, Corp. typically manufactures 1,200,000 moisture gauges per year. Variable costs were $21,000,000 and fixed cost were $1,800,000. Management predicts 1,300,000 gauges will be sold at $50.00 each in the upcoming year. However, the company received special order to produce 300,000 gauges to be sold at a 25% discount off the regular price. The company has the capacity to produce the extra gauges and fixed costs would not increase due to the special order.

Required: Should Crop Monitor Corp accept the order? If they do, what will be the change in income?
Answer #1
Special order price 37.50 =50*(1-25%)
Incremental sales revenue 11250000 =300000*37.50
Less: Incremental variable costs 5250000 =21000000/1200000*300000
Incremental income 6000000
Yes, Crop Monitor Corp should accept the order
Income will increase by $6000000